### v1

@@ 2 Architecture and Core Components
Variables are represented as consecutive indices starting from 0. (In)equations are represented as Equation objects. Piecewise-linear constraints are represented by objects of classes that inherit from the PiecewiseLinearConstraint abstract class. The abstract class defines the key interface methods that are implemented in each sub-class. This way, all piecewise-linear constraints are handled uniformly in the back end. Similarly, each other type of non-linear constraint is implemented as a sub-class of the new NonlinearConstraint abstract class. Initially, Marabou only supported the ReLU and Max constraints. In Marabou 2.0, over ten types of non-linear constraints (listed in Appendix  0.A.1) are supported.
@@ 2.1 Engine
Over the past few years, numerous papers (e.g.,  [65, 56, 77], inter alia) have proposed abstract interpretation techniques that rely on network-level reasoning (e.g., propagating the input bounds layer by layer to tighten output bounds). These analyses can be viewed as a stand-alone, incomplete DNN verification procedure, or as in-processing bound tightening passes for the SMT Solver. Marabou 2.0 features a brand new NetworkLevelReasoner class that supports this type of analysis. The class maintains the neural network topology as a directed acyclic graph, where each node is a Layer object. The Layer class records key information such as weights, source layers, and mappings between neuron indices and variable indices. Currently, seven different analyses are implemented: [i] 1. interval bound propagation [24]; 2. symbolic bound propagation [65]; 3. DeepPoly/CROWN analysis [56, 77]; 4. LP-based bound tightening [60]; 5. Forward-backward analysis [69]; 6. MILP-based bound tightening [60]; and 7. iterative propagation [67]. Analyses 2–7 are implemented in a parallelizable manner, and analyses 4–7 require calls to an external LP solver. By default, the DeepPoly/CROWN analysis is performed. The Network-level Reasoner tightly interleaves with the SMT Solver: the network-level reasoning is executed any time a new search state is reached (with the most up-to-date variable bounds), and the derived bound tightenings are immediately fed back to the search procedure.
@@ 2.4 Front End
Marabou provides interfaces to prepare input queries and invoke the back-end solver in multiple ways. The Marabou executable can be run on the command line, taking in network/property/query files in supported formats. The Python and C++ APIs support this functionality as well, but also contain methods to add arbitrary linear and (supported) non-linear constraints. In addition, a layer on top of the Python API was added to Marabou 2.0 which allows users to define constraints in a more Pythonic manner, resulting in more succinct code. An example is given in Appendix. 0.A.2.
@@
2.5 License and Installation
Marabou is available under the permissive modified BSD open-source license, and runs on Linux and macOS machines. The tool can be built from scratch using CMake. Marabou is now also available on The Python Package Index (PyPI) and can be installed through pip.
@@ 3 Highlighted Features and Applications
In terms of performance, Marabou is on par with state-of-the-art verification tools. In the latest VNN-COMP [11], Marabou won the second place overall, and scored the highest among all CPU-based verifiers. We summarize the main results in Appendix 0.B. In this section, we focus on the usability aspect of Marabou, and highlight some of its recent applications — as well as the features that make them possible. We believe this diverse set of use cases (as well as relevant scripts to be released in the artifact) serve as valuable examples, which will inspire new ways to apply the solver. A runtime evaluation of Marabou 2.0 against an early version appears in Section 4.
@@
A well-studied set of benchmarks in DNN verification derives from an implementation of the ACAS-Xu airborne system for collision avoidance [38]. Using Marabou, we were able to produce certificates of unsatisfiability for these benchmarks for the first time. Marabou was able to produce certificates for 113 out of the 180 tested benchmarks, with only mild overhead incurred by proof generation and certification. The proof certificates contained over 1.46 million proof-tree leaves, of which more than $99\%9$ were certified by the native proof checker, while the remaining were certified by a trusted SMT solver. Additional details are provided in [36].
@@ 4 Runtime Evaluation
Figure 2 compares the runtime of the two Marabou versions on all the benchmarks with a 1 hour CPU timeout. Each configuration was given 1 core and 8GB of memory. Note that Marabou 2.0 was not configured with external solvers in this experiment. We see that Marabou 2.0 is significantly more efficient for a vast majority of the instances. Upon closer examination, an at-least 2×2$\times$2 × speed-up is achieved on 417 instances and an at-least 10×10$\times$10 × speed-up is achieved on 253 instances. Moreover, Marabou 2.0 is also significantly more memory efficient, with a median peak usage of 59MB (versus 604MB with the old version). Solvers’performance on individual benchmarks is reported in Appendix 0.C.2.
@@ Appendix 0.A Architecture and Core Components
0.A.2 Pythonic Interface Example
We provide an example to illustrate the newly introduced Pythonic interface. In the example, we check whether the first output of a network (stored in the ONNX format) can be less than or equal to half of its second output, when the first input is greater than or equal to 0.10.10.1. The base Python API to implement this is shown in (a), and (b) exhibits the “Pythonic” API.
Q = Marabou.read_onnx("model.onnx")
X, Y = Q.inputVars[0], Q.outputVars[0]
Q.setLowerBound(X[0], 0.1)
Q.addInequality([Y[0], Y[1]], [1, -0.5], 0)
Q.solve()
((a)) The base Python API.
Q = Marabou.read_onnx("model.onnx")
X, Y = Q.inputVars[0], Q.outputVars[0]
Q.addConstraint(Var(X[0]) >= 0.1)
Q.addConstraint(Var(Y[0]) <= 0.5 * Var(Y[1]))
Q.solve()
((b)) The “Pythonic” API.

### v2

@@ 2 Architecture and Core Components
Variables are represented as consecutive indices starting from 0. (In)equations are represented as Equation objects. Piecewise-linear constraints are represented by objects of classes that inherit from the PiecewiseLinearConstraint abstract class. The abstract class defines the key interface methods that are implemented in each sub-class. This way, all piecewise-linear constraints are handled uniformly in the back end. Similarly, each other type of non-linear constraint is implemented as a sub-class of the new NonlinearConstraint abstract class. Initially, Marabou only supported the ReLU and Max constraints. In Marabou 2.0, over ten types of non-linear constraints (listed in Appendix 0.A.1) are supported.
@@ 2.1 Engine
Over the past few years, numerous papers (e.g.,  [66, 57, 80, 50, 78], inter alia) have proposed abstract interpretation techniques that rely on network-level reasoning (e.g., propagating the input bounds layer by layer to tighten output bounds). These analyses can be viewed as a stand-alone, incomplete DNN verification procedure, or as in-processing bound tightening passes for the SMT Solver. Marabou 2.0 features a brand new NetworkLevelReasoner class that supports this type of analysis. The class maintains the neural network topology as a directed acyclic graph, where each node is a Layer object. The Layer class records key information such as weights, source layers, and mappings between neuron indices and variable indices. Currently, seven different analyses are implemented: [i] 1. interval bound propagation [24]; 2. symbolic bound propagation [66]; 3. DeepPoly/CROWN analysis [57, 80]; 4. LP-based bound tightening [61]; 5. Forward-backward analysis [70]; 6. MILP-based bound tightening [61]; and 7. iterative propagation [68]. Analyses 2–7 are implemented in a parallelizable manner, and analyses 4–7 require calls to an external LP solver. By default, the DeepPoly/CROWN analysis is performed. The Network-level Reasoner tightly interleaves with the SMT Solver: the network-level reasoning is executed any time a new search state is reached (with the most up-to-date variable bounds), and the derived bound tightenings are immediately fed back to the search procedure.
@@ 2.4 Front End
Marabou provides interfaces to prepare input queries and invoke the back-end solver in multiple ways. The Marabou executable can be run on the command line, taking in network/property/query files in supported formats. The Python and C++ APIs support this functionality as well, but also contain methods to add arbitrary linear and (supported) non-linear constraints. In addition, a layer on top of the Python API was added to Marabou 2.0 which allows users to define constraints in a more Pythonic manner, resulting in more succinct code. For example, suppose one wants to check whether the first output of a network (stored in the ONNX format) can be less than or equal to half of its second output, when the first input is greater than or equal to $0.1$. Figure 2(a) shows how to perform this check with the base Python API, while Figure 2(b) exhibits the “Pythonic” API.
Q = Marabou.read_onnx("model.onnx")
X, Y = Q.inputVars[0], Q.outputVars[0]
Q.setLowerBound(X[0], 0.1)
Q.addInequality([Y[0], Y[1]], [1, -0.5], 0)
Q.solve()
((a)) The base Python API
Q = Marabou.read_onnx("model.onnx")
X, Y = Q.inputVars[0], Q.outputVars[0]
Q.addConstraint(Var(X[0]) >= 0.1)
Q.addConstraint(Var(Y[0]) <= 0.5 * Var(Y[1]))
Q.solve()
((b)) The “Pythonic” API
Figure 2: Two ways to define the same verification query through the Python API.
@@
2.5 Availability, License, and Installation
Marabou is available under the permissive modified BSD open-source license, and runs on Linux and macOS machines. The tool can be built from scratch using CMake. Marabou is now also available on The Python Package Index (PyPI) and can be installed through pip. The latest version of Marabou is available at: https://github.com/NeuralNetworkVerification/Marabou. The artifact associated with this tool description is archived on Zenodo [71].
@@ 3 Highlighted Features and Applications
In terms of performance, Marabou is on par with state-of-the-art verification tools. In the latest VNN-COMP [11], Marabou won the second place overall, and scored the highest among all CPU-based verifiers. We summarize the main results in Appendix 0.B. In this section, we focus on the usability aspect of Marabou, and highlight some of its recent applications — as well as the features that make them possible. We believe this diverse set of use cases (as well as the relevant scripts in the artifact [71]) serve as valuable examples, which will inspire new ways to apply the solver. A runtime evaluation of Marabou 2.0 against an early version appears in Section 4.
@@
A well-studied set of benchmarks in DNN verification derives from an implementation of the ACAS-Xu airborne system for collision avoidance [38]. Using Marabou, we were able to produce certificates of unsatisfiability for these benchmarks for the first time. Marabou was able to produce certificates for 113 out of the 180 tested benchmarks, with only mild overhead incurred by proof generation and certification. The proof certificates contained over 1.46 million proof-tree leaves, of which more than $99.99\%$ were certified by the native proof checker, while the remaining were certified by a trusted SMT solver. Additional details are provided in [36].
@@ 4 Runtime Evaluation
Figure 3 compares the runtime of the two Marabou versions on all the benchmarks with a 1 hour CPU timeout. Each configuration was given 1 core and 8GB of memory. Note that Marabou 2.0 was not configured with external solvers in this experiment. We see that Marabou 2.0 is significantly more efficient for a vast majority of the instances. Upon closer examination, an at-least 2×2$\times$2 × speed-up is achieved on 428 instances and an at-least 10×10$\times$10 × speed-up is achieved on 263 instances. Moreover, Marabou 2.0 is also significantly more memory efficient, with a median peak usage of 57MB (versus 604MB with the old version). Solvers’performance on individual benchmarks is reported in Appendix 0.C.2.